OpenStudy (anonymous):

Solve the equation 15(x - 3) = 12x - 43. A. 2/3 B. 3/2 C. -40/3 D. 3

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

First, distribute the 15 to x and -3

OpenStudy (anonymous):

15x and -45

OpenStudy (anonymous):

add 45 to both sides

OpenStudy (anonymous):

what do you mean can you draw it out ?

OpenStudy (anonymous):

15(x - 3) = 12x - 43 <=>15x-45=12x-43<=>15x-12x=45-43<=>3x=2<=>x=2/3 AAAAAAAAAAA

OpenStudy (anonymous):